### v1

# SLANG: New Concept Comprehension of Large Language Models

Abstract
The dynamic nature of language, particularly evident in the realm of slang and memes on the Internet, poses serious challenges to the adaptability of large language models (LLMs). Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities. This research addresses the critical need to bridge this gap, aiming to enhance LLMs’ comprehension of evolving new concepts on the internet, without the high cost and impracticality of continual retraining. To address this issue, we propose a new benchmark SLANG to assess LLMs’ proficiency in comprehending emerging linguistic trends and a baseline approach FOCUS, which uses causal inference to enhance LLMs to understand new phrases and usage patterns. This approach involves scrutinizing real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their intended meanings. The empirical analysis shows that our causal inference-based approach outperforms the traditional models in terms of precision and relevance in the interpretation of Internet slang and memes. [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
SLANG: New Concept Comprehension of Large Language Models
[…]
1 Introduction
The evolution of human language is the history of human’s innovation and adaptation. From the earliest forms of communication to the complex linguistic systems of today, language has continuously transformed, reflecting, and shaping the human experience. Recently, the language evolution has been accelerated by the developing communication technology, particularly the Internet, which has introduced new dimensions to linguistic evolution (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Figure 1: Comparative analysis of LLMs’ understanding of new phrases using CoT Wei et al. (2022) and FOCUS methods. The left side demonstrates the limited understanding through the CoT approach, focusing on the literal interpretation. In contrast, the right side using the FOCUS method shows the model’s enhanced capability to grasp metaphors and deeper meanings.
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and knowledge emerge at an unprecedented pace (Sun et al., 2021). The crux of our research lies in exploring the potential of LLMs to adapt to and understand these linguistic transformations and new knowledge without resorting to continual retraining or fine-tuning with updated datasets or retrieving external data.
Note that many popular LLMs are often mainly black-box models. They may make decisions based on superficial patterns rather than justified reasons, which in turn can affect the model’s ability to understand and follow human instructions within context (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023). For example, as shown in Figure 1, while CoT (Chain-of-Thought) Wei et al. (2022) provides a straightforward interpretation of the phrase "The sunset is beautiful, isn’t it?", our method recognizes the underlying metaphorical meaning, identifying it as an acknowledgment of an ending, such as a concluding phase in life or a relationship, in a nuanced conversational exchange. This recognition calls for a more nuanced approach in the development and evaluation of LLMs, which considers not only their performance metrics, but also the underlying principles guiding their decision-making processes.
In this study, we propose FOCUS (Factual cOntext CaUsal AnalysiS), a new methodology centered on causal inference, which can significantly enhance the language model’s capability to understand new phrases. This approach is grounded in the belief that understanding causal relationships in language usage, especially in evolving linguistic contexts, leads to more accurate and adaptable models. Our method involves constructing new datasets from UrbanDictionary Urban Dictionary LLC , comprising both factual and counterfactual instances. These datasets provide diverse linguistic contexts, ranging from common usage to generative scenarios, aiding in the comprehensive understanding of new phrases and concepts.
In addition to FOCUS, we introduce SLANG (Similarity of Lexical Analysis aNd Grasp), a new benchmark designed to evaluate language models’ adaptability to linguistic evolution and vocabulary changes. The SLANG benchmark focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating our FOCUS methodology.
2 SLANG
In response to the rapid evolution of the Internet language, we introduce the SLANG benchmark, a tool derived from UrbanDictionary contributions. SLANG aims to evaluate the proficiency of LLMs in interpreting the ever-changing landscape of user-generated content. This benchmark is unique in its division into factual and counterfactual datasets, each serving a distinct purpose in understanding the adaptability of LLMs. The following subsections explore the definition of the SLANG problem, detail the dataset, and discuss the metrics used for evaluation. This structure ensures a comprehensive assessment of LLMs’ capabilities in navigating the complexities of modern linguistic trends.
2.1 Problem Definition
Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs. Our journey unfolds with the task of capturing the essence of this linguistic fluidity, mathematically expressed as the challenge of estimating $P(Y|W,X;\mathcal{M})$. This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
2.2 Dataset Constructions
Data Extraction
In the construction of the dataset, our methodology involved the systematic extraction of a substantial number of concepts from UrbanDictionary Urban Dictionary LLC , a widely recognized and reliable online repository of Internet slang and colloquial terms. UrbanDictionary stands out for its user-driven content, reflecting current language trends and the evolving nature of the internet lexicon. This platform serves as both a repository and a dynamic community, continuously enriching the evolving landscape of internet language with new concepts. Specifically, our approach involved selecting concepts that were added after a predetermined date, a strategy that was integral to ensuring the relevance and recency of the dataset. The subsequent steps undertaken for the development of this dataset were as follows. Our dataset’s construction commenced with UrbanDictionary as the foundational source. We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes). This approach ensured that our dataset was grounded in authentic user-driven language trends. Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
Figure 2: Comparative histograms illustrating the distribution of upvotes (left) and absolute upvotes (right) across dataset entries. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning.
Content Filtering
The content filtering of our dataset consists of several critical steps, with the foremost being the assurance of content novelty. Millions of concepts on UrbanDictionary have led to a significant proportion of phrases becoming commonplace in everyday communication. This prevalence implies that many of these phrases might already be included in GPT’s training corpus, as seen in the case of widely recognized terms like "ghosting" or "flex", which have transitioned from internet slang to general usage. As such, predicting the meanings of these phrases using LLMs would be redundant. Therefore, our primary objective was to ensure that the phrases and their definitions in our dataset were unfamiliar to the LLMs. In alignment with this objective, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
An essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
We then rigorously employ gpt-4-0613 for an exhaustive filtration process, querying each term repeatedly to ensure that they were not memrozied by the GPT, thereby ensuring the unique and contemporary nature of our dataset.
Additionally, we rigorously filtered inappropriate content to uphold academic integrity, thoroughly examining each entry for explicit, violent, or harmful material to ensure the dataset’s appropriateness and quality.
Factual Dataset
Acknowledging the nonformal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. However, considering that language models might output synonymous interpretations with varied wording, we augmented each transformed explanation with four additional variants incorporating synonym replacement. This approach was designed to capture the diversity of potential language model outputs. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
Counterfactual Dataset
Another core of our methodology lies in the creation of a counterfactual context. We aim to evaluate the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding. This approach is characterized by the preservation of original meanings from the factual dataset, while simultaneously introducing entirely new phrases, thereby shifting the linguistic form while maintaining the semantic substance.
This refined strategy requires the innovation of novel phrases that are linguistically distinct from their real-world counterparts, yet congruent with the retained meanings. The challenge lies in crafting these new phrases to be unique in structure, yet resonant with the established meanings, thus presenting a different verbal manifestation of the same conceptual core.
The process involves a creative and analytically rigorous approach, where the rich contextual foundation of the original dataset serves as the semantic basis for the generation of new phrases. These phrases are designed to encapsulate the same ideas as the original meanings, but articulated through a different linguistic prism. Achieving a balance between inventive language use and adherence to logical linguistic norms is crucial to ensuring that new phrases are plausible and coherent within established linguistic frameworks.
Upon generation, these new phrases are systematically paired with their corresponding original meanings, creating a dataset that juxtaposes novel linguistic expressions against established semantic structures. This pairing enables a multifaceted analysis of how LLMs process and comprehend language, specifically assessing their capability to disentangle linguistic form from semantic substance. This methodology offers insight into the nuanced capabilities of LLMs in interpreting language beyond mere recognition, exploring the realms of semantics and contextual linguistics.
2.3 Metrics
Given the nature of this task as a question-answering (QA) problem, traditional QA metrics like F1 score, Recall, and Precision were employed Yang et al. (2018). Moreover, we introduced BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) as metrics to explore the capacity of the listed methods to generate higher quality interpretations with stricter similarity requirements. Again, considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) is obtained by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
3 FOCUS
Figure 3: Structural Causal Model (SCM) of large language models for interpreting new phrases, excluding confounders. The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge. Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
This section outlines the novel methodology employed in our research to enhance the adaptability of LLMs in understanding the evolving human language. Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
3.1 Causal Analysis
Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables. We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context. IIn this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation. These causal links, free from confounders, capture the logical chain from input to interpretation. This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
Figure 4: SCM analysis in FOCUS methodology. This figure presents the complex process of causal inference in the understanding of new phrases, highlighting how the FOCUS approach systematically analyzes and interprets intricate language patterns, emphasizing the causal links between linguistic elements and their interpretive outcomes.
3.1.1 Causal Analysis of Entity
Initially, we reference a typic SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
Figure 5: The four-stage pipeline of FOCUS.
3.1.2 Analysis with Linguistic Factors
In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables. The model includes direct paths $X\rightarrow Y$ and $W\rightarrow Y$, indicating the immediate influence of context and phrase on the interpretation. The backdoor path $E\leftarrow X\rightarrow W$ and the indirect path $X\rightarrow E\rightarrow W$ demonstrate the mediated effects. The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context. This configuration captures the causal mechanisms underlying natural language understanding, emphasizing the direct, indirect, and mediated relationships among the variables. The equation for the causal effect in this context is as follows:
$P(Y=y|X=x)$
$=$ $\sum_{e,r}P(Y=y|X=x,W=w,E=e,R=r)$
$P(W=w|X=x,E=e,R=r)$
$P(E=e)P(R=r)$
In this formula, $P(Y=y|X=x)$ represents the probability of the outcome variable $Y$ being a particular value $y$, given the context $X$ is set to $x$. The summation over $e$ and $r$ encompasses all possible combinations of the values of the entities $E$ and linguistic factors $R$. For each combination, the formula calculates the conditional probability of $Y=y$ given $X=x$, the specific phrase $W=w$, the entity $E=e$, and the linguistic factor $R=r$. This conditional probability is further modulated by the probabilities $P(W=w|X=x,E=e,R=r)$, $P(E=e)$, and $P(R=r)$. These terms represent the likelihood of observing the phrase $W=w$ conditioned on the context, entities, and linguistic factors, as well as the inherent probabilities of the entities $E=e$ and the linguistic factors $R=r$. This comprehensive approach allows for a nuanced understanding of how context, entities, and linguistic factors collectively influence the interpretation $Y$.
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
CoT Wei et al. (2022) 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
CauView Wang et al. (2023) 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
FOCUS (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
Table 1: Performance results on the factual dataset.
3.2 Context-based Causal Intervention
Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors. To eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (d)). This alteration aids in isolating the effect of $W$ while excluding entity-specific biases. Further, to concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$. This shift accentuates the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner. The updated causal effect formula in the SCM is thus:
$P(Y=y|do(X=x))$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|do(X=x),\tilde{E}=\tilde{e},R=r)$
$P(\tilde{E}=\tilde{e})P(R=r)$
$=$ $\sum_{\tilde{e},r}P(Y=y|X=x,\tilde{W}=\tilde{w},\tilde{E}=\tilde{e},R=r)$
$P(\tilde{W}=\tilde{w}|\tilde{E}=\tilde{e},R=r)P(\tilde{E}=\tilde{e})P(R=r)$
This revised approach and formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention. This approach ensures a more robust and bias-free interpretation within the SCM framework. This equation accounts for the altered relationships in the SCM after $do$-operation. The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context. The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language. To achieve this objective, we propose a four-stage method, here is a description of each stage and its principles:
Direct Inquiry (DI)
The DI stage is the foundational step in our FOCUS methodology. It primarily involves inputting the usage example (context) $X$ and the phrase $W$ directly into the language model. This stage is represented as $Y_{DI}=\mathcal{M}(X,W,E)$, where $\mathcal{M}$ is the function that defines the processing of the language model. The aim here is to assess the direct effect of the context $X$ on the inferred meaning $Y$ in a straightforward and intervention-free manner. This allows the model to naturally derive the meaning within the given context. Special emphasis is placed on maintaining the integrity of all confounders, ensuring that the phrase’s morphology, literal meaning, and context-based interpretation are coherently analyzed. The DI stage sets the baseline for understanding the meaning of the phrase before any further analytical layers are added in the subsequent stages.
Masked Entity Analysis (MEA)
The MEA stage is a crucial component of our FOCUS methodology, designed to mitigate the impact of the phrase’s literal meaning on its interpretation. In this stage, the phrase $W$ is masked within the context $X$, leading to an evaluation of the meaning $Y$ in the absence of direct influence from $W$. This process is denoted as $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, where $\mathcal{M}$ represents the language model. The primary objective is to understand the meaning based on context, prompting the model to propose synonyms or near-synonyms that could feasibly replace the masked phrase. The MEA stage allows for a nuanced analysis of how the interpretation $Y$ is influenced when $W$ is obscured, yet still impacted by the entities $E$. This stage is instrumental in mitigating the literal interpretation bias of $W$, focusing instead on deriving meaning from the broader linguistic factors present in $X$ and $E$.
Entity Replacement Inquiry (ERI)
The ERI stage plays a pivotal role in our FOCUS methodology, where we explore the impact of varying entity relationships on the interpretation of phrases to delve deeper into context-based interpretation. In this stage, we implement an entity replacement strategy combined with a dropout rate, selectively altering entities within the context $X$. This process, denoted as $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$, involves generating multiple context sentences with replaced entities, while a dropout rate determines the extent of entity replacement. The addition of this dropout strategy is not only to introduce randomness but also to prevent the emergence of new entity biases that could arise from completely replacing all entities. By maintaining a mix of original and altered entities, we ensure the robustness of understanding while mitigating the risk of introducing new biases. The language model $\mathcal{M}$ then conducts multiple inferences on these varied contexts to synthesize a cohesive interpretation. This approach enhances the robustness of the model by considering a diverse range of entity interactions. The ERI stage is fundamental in understanding the nuanced effects of entity dynamics within different contexts, providing a more comprehensive causal analysis of phrase interpretation.
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
1 2 L
GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
CoT Wei et al. (2022) 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
CauView Wang et al. (2023) 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
FOCUS (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
Table 2: Performance results on the counterfactual dataset.
Final Synthesis (FS)
The FS stage is the integrative phase of the FOCUS methodology, symbolized as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$. This stage synthesizes insights from Direct Inquiry (DI), Masked Entity Analysis (MEA), and Entity Replacement Inquiry (ERI). It evaluates the interplay between direct, contextually-driven, and variable entity interpretations to forge a comprehensive understanding of the phrase $W$ within context $X$ and its relation to entities $E$. In FS, we assess causal relationships, reconciling the varied interpretations and confounding factors identified in previous stages. This synthesis refines our understanding of language model interpretations, emphasizing contextual richness and linguistic nuance. The FS stage thus concludes the FOCUS methodology, offering a multi-dimensional perspective on language model analytics.
4 Experiments
4.1 Setup
Large Language Models
In our experimental setup, we utilized a subset of 408 samples, filtered from our initial dataset of 7220 concepts, focusing on evaluations over GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613). These models were selected for their advanced capabilities and the relevance of their knowledge bases, which have a cut-off date prior to April 2021. As we discussed in 2.2, this cut-off is crucial to ensure the novelty and relevancy of our dataset, as it aligns with the temporal frame of our selected UrbanDictionary concepts.
Baselines
Baselines comprised direct inquiry (Direct) Ouyang et al. (2022), Chain-of-Thought (CoT) Wei et al. (2022), CauView Wang et al. (2023), and our FOCUS approaches. In each case, the language model output was parsed and compared with the ground truth. We implement the CauView method in our experiment design by using a two-stage prompt inquiry, due to its lack of a direct inquiry step.
FOCUS
The outputs from the first three stages were processed by a parser that retained only the final conclusions, filtering out the model’s “reasoning" process. In the fourth stage, the language model was briefed about the concise procedures and objectives of the previous three methods. This briefing included potential key information each method might yield and the risks of information omission. Additionally, to control for syntactic variability and minimize error, the prompts for all baselines, including FOCUS, incorporated the same three sample entries for few-shot in-context learning. This ensured that the language model’s output syntax closely resembled that of the dataset.
Experiment F1 Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
w/o MEA 0.4366 0.4380 0.4484 0.2766 0.4586 0.2059 0.3556 0.8821 0.8014 82.0
w/o ERI 0.4283 0.4300 0.4371 0.2814 0.4593 0.2117 0.3547 0.9021 0.8092 84.0
Table 3: Results of ablation experiments on GPT-4.
4.2 Results
Factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We observed that the FOCUS consistently outperformed other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieved the highest F1 score of 0.4292, Precision of 0.4153, and Recall of 0.4524, along with the best BLUE, ROUGE, Similarity, SimCSE scores, and Accuracy at 84.5%. A similar trend was noted with the GPT-4 model. Here, FOCUS again led the metrics, notably achieving an F1 score of 0.4446, Precision of 0.4280, Recall of 0.4714, and an impressive Accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies. The comparative analysis clearly illustrates that our FOCUS approach excels in understanding the nuanced and evolving nature of internet language, thereby marking a significant advancement in the field of natural language processing.
Counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again excelled, outperforming others in metrics with GPT-4, achieving an F1 score of 0.4532, Precision of 0.4598, Recall of 0.4551, and an Accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. Comparatively, on the factual dataset, while FOCUS led with substantial margins, the gap in performance was more pronounced in the counterfactual dataset. This discrepancy underscores the enhanced complexity in interpreting altered linguistic contexts, a challenge adeptly met by FOCUS. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
4.3 Ablation Study
Our ablation study, focusing on the FOCUS methodology’s components MEA and ERI, utilized GPT-4 on the factual dataset. As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
MEA (Masked Entity Analysis)
MEA, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance. Its exclusion (w/o MEA) reduced the F1 score to 0.4366 from 0.4446. This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
ERI (Entity Replacement Inquiry)
The ERI process, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact. Removing ERI (w/o ERI) decreased the F1 score to 0.4283. ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
While incorporating either MEA or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results. This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
5 Related Work
5.1 Knowledge Update Methods
LLMs enhance knowledge through Parameter-Efficient Fine-Tuning (PEFT) methods like task-specific parameter addition Houlsby et al. (2019) and low-rank adaptation (LoRA) Hu et al. (2021); Pfeiffer et al. (2020). Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al. (2021). BitFit Ben-Zaken et al. (2021), while simplifying fine-tuning, relies heavily on dataset quality. Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods. Retrieval-Augmented Generation Lewis et al. (2020) and In-context learning-based knowledge editing Zhong et al. (2023) offer dynamic integration of external information, focusing more on fact retrieval than on enhancing deeper understanding.
5.2 Entity Bias and Shortcuts in LLMs
Entity bias Peng et al. (2020); Longpre et al. (2021); Wang et al. (2022, 2023) and shortcuts Du et al. (2021); Saparov and He (2023) in LLMs lead to oversimplified language processing, relying on specific entities or dataset-driven patterns. Entity bias skews model predictions towards certain entities, while shortcuts encompass simplified heuristics, focusing on identifiable features or aspects of the input Du et al. (2022). These patterns limit the models’ understanding and generation of nuanced language, affecting generalization and robustness.
5.3 Causal Intervention Solutions
Causal interventions for debiasing and mitigating shortcuts have gained prominence. Tian et al. (2022) and Zhou et al. (2023a) focus on causal inference and invariant learning for debiasing. CausaLM Feder et al. (2021) provides causal-based model explanations, addressing previous tools’ limitations. Counterfactual methods for debiasing Chen et al. (2023) and eliminating shortcuts Wen et al. (2022) have shown promise, though limitations remain in targeting white-box models and retraining requirements. Our research addresses these limitations through contextual learning and causal inference for bias and shortcut elimination without retraining.
6 Conclusion
In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs). Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends. Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new phrases and usage patterns, going beyond traditional models in precision and relevance. Our analysis revealed that the causal inference-based approach of FOCUS significantly outperforms traditional models in interpreting internet slang and memes. This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts. The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our FOCUS method, to adapt to the rapid evolution of online language. This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
7 Acknowledge
This paper is partially supported by the National Science Foundation of China under Grant No.U1911401, U21B2046 6237075198.

### v4

@@
 Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities.
-This research addresses the critical need to bridge this gap, aiming to enhance LLMs’ comprehension of evolving new concepts on the internet, without the high cost and impracticality of continual retraining.
-To address this issue, we propose a new benchmark SLANG to assess LLMs’ proficiency in comprehending emerging linguistic trends and a baseline approach FOCUS, which uses causal inference to enhance LLMs to understand new phrases and usage patterns.
-This approach involves scrutinizing real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their intended meanings.
-The empirical analysis shows that our causal inference-based approach outperforms the traditional models in terms of precision and relevance in the interpretation of Internet slang and memes.
+This research aims to bridge this gap by enhancing LLMs’ comprehension of the evolving new concepts on the Internet, without the high cost of continual retraining.
+In pursuit of this goal, we introduce SLANG, a benchmark designed to autonomously integrate novel data and assess LLMs’ ability to comprehend emerging concepts, alongside FOCUS, an approach uses causal inference to enhance LLMs to understand new phrases and their colloquial context.
+Our benchmark and approach involves understanding real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their meanings.
+The empirical analysis shows that our causal inference-based approach outperforms the baseline methods in terms of precision and relevance in the comprehension of Internet slang and memes.
 [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
@@
 1 Introduction
-The evolution of human language is the history of human’s innovation and adaptation.
-From the earliest forms of communication to the complex linguistic systems of today, language has continuously transformed, reflecting, and shaping the human experience.
-Recently, the language evolution has been accelerated by the developing communication technology, particularly the Internet, which has introduced new dimensions to linguistic evolution (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016).
+Recently, the language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016).
 These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
+Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language.
+This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021).
+Hence, it is essential for LLMs to to understand linguistic shifts and new concepts without constant updates or external data.
 Figure 1: Comparative analysis of LLMs’ understanding of new phrases using CoT Wei et al.
@@
 In contrast, the right side using the FOCUS method shows the model’s enhanced capability to grasp metaphors and deeper meanings.
-Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language.
-This limitation is particularly pronounced in the context of digital communication, where new forms of expression and knowledge emerge at an unprecedented pace (Sun et al., 2021).
-The crux of our research lies in exploring the potential of LLMs to adapt to and understand these linguistic transformations and new knowledge without resorting to continual retraining or fine-tuning with updated datasets or retrieving external data.
-Note that many popular LLMs are often mainly black-box models.
-They may make decisions based on superficial patterns rather than justified reasons, which in turn can affect the model’s ability to understand and follow human instructions within context (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023).
-For example, as shown in Figure 1, while CoT (Chain-of-Thought) Wei et al.
-(2022) provides a straightforward interpretation of the phrase "The sunset is beautiful, isn’t it?", our method recognizes the underlying metaphorical meaning, identifying it as an acknowledgment of an ending, such as a concluding phase in life or a relationship, in a nuanced conversational exchange.
-This recognition calls for a more nuanced approach in the development and evaluation of LLMs, which considers not only their performance metrics, but also the underlying principles guiding their decision-making processes.
-In this study, we propose FOCUS (Factual cOntext CaUsal AnalysiS), a new methodology centered on causal inference, which can significantly enhance the language model’s capability to understand new phrases.
-This approach is grounded in the belief that understanding causal relationships in language usage, especially in evolving linguistic contexts, leads to more accurate and adaptable models.
-Our method involves constructing new datasets from UrbanDictionary Urban Dictionary LLC , comprising both factual and counterfactual instances.
-These datasets provide diverse linguistic contexts, ranging from common usage to generative scenarios, aiding in the comprehensive understanding of new phrases and concepts.
-In addition to FOCUS, we introduce SLANG (Similarity of Lexical Analysis aNd Grasp), a new benchmark designed to evaluate language models’ adaptability to linguistic evolution and vocabulary changes.
-The SLANG benchmark focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating our FOCUS methodology.
+Moreover, LLMs often make decisions based on superficial patterns rather than justified reasons.
+This can hinder their ability to accurately interpret and follow human instructions, as highlighted in several studies (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023).
+For example, as depicted in Figure 1, the Chain-of-Thought (CoT) prompting Wei et al.
+(2022) simply interprets the phrase The sunset is beautiful, isn’t it?
+and misses the deeper, metaphorical meaning, which could represent an acknowledgment of an ending, like concluding a life phase or a relationship, in a complex conversation.
+This situation underscores the importance of enhancing and evaluating LLMs in a way that goes beyond their performance metrics.
+It’s essential to consider the fundamental principles that guide their decision-making processes.
+Therefore, we propose SLANG (Similarity of Lexical Analysis aNd Grasp), a benchmark to assess language models’ adaptability to linguistic shifts, and FOCUS (Factual cOntext CaUsal analysiS), an approach based on causal inference for enhancing comprehension of new concepts.
+The SLANG benchmark, developed from UrbanDictionary Urban Dictionary LLC , focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating the capability of LLMs in grasping new concepts.
+We select recent entries after a specific cutoff date and filter out phrases already likely in LLM training data.
+We utilize user-generated ratings (upvotes and downvotes) to refine the dataset, ensuring its quality and comprehensiveness.
+The dataset is then standardized into a formal dictionary format, simplifying explanations and examples for universal understanding while retaining original meanings.
+This preprocessing approach ensures SLANG effectively evaluates LLMs’ adaptability to linguistic shifts.
+FOCUS employs causal inference to enhance models’ comprehension of new concepts within evolving linguistic contexts.
+By analyzing causal relationships in language, FOCUS advances models’ predictive capabilities beyond traditional correlation-based learning.
+This method allows for a nuanced grasp of language dynamics, improving models’ adaptability and effectiveness in applications requiring deep understanding of language use.
+FOCUS significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability.
+With GPT-4, FOCUS achieved an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, and recall of 0.4551, alongside an accuracy of 84.9%, outperforming previous methods in comprehension and adaptability.
+The codes for the SLANG and FOCUS toolboxes are open-sourced, contributing to the community’s resources for advancing language model development.
 2 SLANG
-In response to the rapid evolution of the Internet language, we introduce the SLANG benchmark, a tool derived from UrbanDictionary contributions.
-SLANG aims to evaluate the proficiency of LLMs in interpreting the ever-changing landscape of user-generated content.
-This benchmark is unique in its division into factual and counterfactual datasets, each serving a distinct purpose in understanding the adaptability of LLMs.
-The following subsections explore the definition of the SLANG problem, detail the dataset, and discuss the metrics used for evaluation.
-This structure ensures a comprehensive assessment of LLMs’ capabilities in navigating the complexities of modern linguistic trends.
-2.1 Problem Definition
+We introduce SLANG in response to rapidly evolving language.
+SLANG benchmark evaluates LLMs’ capability to interpret the dynamic landscape of user-generated new concepts.
+It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability.
+We detail SLANG’s problem definition, dataset construction, and evaluation metrics in the following subsections.
+2.1 Problem
 Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs.
@@
 This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
-2.2 Dataset Constructions
-Data Extraction
-In the construction of the dataset, our methodology involved the systematic extraction of a substantial number of concepts from UrbanDictionary Urban Dictionary LLC , a widely recognized and reliable online repository of Internet slang and colloquial terms.
-UrbanDictionary stands out for its user-driven content, reflecting current language trends and the evolving nature of the internet lexicon.
-This platform serves as both a repository and a dynamic community, continuously enriching the evolving landscape of internet language with new concepts.
-Specifically, our approach involved selecting concepts that were added after a predetermined date, a strategy that was integral to ensuring the relevance and recency of the dataset.
-The subsequent steps undertaken for the development of this dataset were as follows.
-Our dataset’s construction commenced with UrbanDictionary as the foundational source.
+2.2 Preprocessing
+Extraction
+Our dataset construction involved extracting numerous concepts from UrbanDictionary, a platform known for user-generated content that reflects current language trends and the evolving internet lexicon, making it a unique, constantly updated repository and dynamic forum for new Internet language concepts.
+Specifically, our approach involved selecting concepts added after a predetermined date to ensure content novelty.
 We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes).
-This approach ensured that our dataset was grounded in authentic user-driven language trends.
 Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
+Filtering
+The content filtering of our dataset consists of several steps, with the foremost being the assurance of content novelty.
+Many UrbanDictionary phrases are now common, potentially pre-included in GPT’s training data.
+Terms like ghosting or flex illustrate this shift to general usage.
+To avoid redundancy in LLM predictions, we focused on ensuring that the phrases and definitions in our dataset were new to the LLMs.
+To solve this, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models.
+To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022.
+This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
+Another essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2.
+Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process.
+This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
+Additionally, we employed gpt-4-0613 for thorough filtration, ensuring terms were not already memorized by GPT, thus maintaining our dataset’s novelty.
+We also removed inappropriate content, preserving academic integrity and ensuring quality.
+Factual dataset
+Acknowledging the informal nature of the original dataset, we transformed the data into a uniform, formal dictionary format.
+This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning.
+The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts.
+We also enriched each explanation with four synonym-based variants, ensuring the capture of diverse potential responses.
+When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches.
+This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
+Counterfactual dataset
+We aim to assess the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding by preserving original meanings from the factual dataset while introducing new phrases.
+This approach maintains the semantic substance but shifts the linguistic form.
+Our strategy involves creating novel phrases that are linguistically distinct yet semantically congruent with their real-world counterparts.
+Crafting these phrases requires a creative and analytically rigorous approach, leveraging the rich contextual foundation of the original dataset.
+These new phrases encapsulate the same ideas as the original meanings but are expressed through a different linguistic perspective.
+Balancing inventive language use with logical linguistic norms is essential to ensure the new phrases are plausible and coherent.
+Upon generation, these new phrases are paired with their corresponding original meanings, creating a dataset that contrasts novel linguistic expressions against established semantic structures.
+This pairing allows for a multifaceted analysis of LLMs’ language processing and comprehension, particularly their ability to separate linguistic form from semantic substance.
+This methodology sheds light on the nuanced capabilities of LLMs in interpreting language beyond mere recognition, delving into semantics and contextual linguistics.
 Figure 2: Comparative histograms illustrating the distribution of upvotes (left) and absolute upvotes (right) across dataset entries.
@@
 The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning.
-Content Filtering
-The content filtering of our dataset consists of several critical steps, with the foremost being the assurance of content novelty.
-Millions of concepts on UrbanDictionary have led to a significant proportion of phrases becoming commonplace in everyday communication.
-This prevalence implies that many of these phrases might already be included in GPT’s training corpus, as seen in the case of widely recognized terms like "ghosting" or "flex", which have transitioned from internet slang to general usage.
-As such, predicting the meanings of these phrases using LLMs would be redundant.
-Therefore, our primary objective was to ensure that the phrases and their definitions in our dataset were unfamiliar to the LLMs.
-In alignment with this objective, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models.
-To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022.
-This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
-An essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2.
-Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process.
-Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes.
-The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning.
-This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
-We then rigorously employ gpt-4-0613 for an exhaustive filtration process, querying each term repeatedly to ensure that they were not memrozied by the GPT, thereby ensuring the unique and contemporary nature of our dataset.
-Additionally, we rigorously filtered inappropriate content to uphold academic integrity, thoroughly examining each entry for explicit, violent, or harmful material to ensure the dataset’s appropriateness and quality.
-Factual Dataset
-Acknowledging the nonformal nature of the original dataset, we transformed the data into a uniform, formal dictionary format.
-This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning.
-The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts.
-However, considering that language models might output synonymous interpretations with varied wording, we augmented each transformed explanation with four additional variants incorporating synonym replacement.
-This approach was designed to capture the diversity of potential language model outputs.
-When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches.
-This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
-Counterfactual Dataset
-Another core of our methodology lies in the creation of a counterfactual context.
-We aim to evaluate the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding.
-This approach is characterized by the preservation of original meanings from the factual dataset, while simultaneously introducing entirely new phrases, thereby shifting the linguistic form while maintaining the semantic substance.
-This refined strategy requires the innovation of novel phrases that are linguistically distinct from their real-world counterparts, yet congruent with the retained meanings.
-The challenge lies in crafting these new phrases to be unique in structure, yet resonant with the established meanings, thus presenting a different verbal manifestation of the same conceptual core.
-The process involves a creative and analytically rigorous approach, where the rich contextual foundation of the original dataset serves as the semantic basis for the generation of new phrases.
-These phrases are designed to encapsulate the same ideas as the original meanings, but articulated through a different linguistic prism.
-Achieving a balance between inventive language use and adherence to logical linguistic norms is crucial to ensuring that new phrases are plausible and coherent within established linguistic frameworks.
-Upon generation, these new phrases are systematically paired with their corresponding original meanings, creating a dataset that juxtaposes novel linguistic expressions against established semantic structures.
-This pairing enables a multifaceted analysis of how LLMs process and comprehend language, specifically assessing their capability to disentangle linguistic form from semantic substance.
-This methodology offers insight into the nuanced capabilities of LLMs in interpreting language beyond mere recognition, exploring the realms of semantics and contextual linguistics.
 2.3 Metrics
-Given the nature of this task as a question-answering (QA) problem, traditional QA metrics like F1 score, Recall, and Precision were employed Yang et al.
-(2018).
-Moreover, we introduced BLEU (3-gram) Papineni et al.
-(2002) and ROUGE Lin (2004) as metrics to explore the capacity of the listed methods to generate higher quality interpretations with stricter similarity requirements.
-Again, considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al.
-(2021), and sentence similarity Team (2023) is obtained by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020).
+For this task, we employed traditional metrics like $\text{F}_{1}$ score, recall, and precision Yang et al.
+(2018), and added BLEU (3-gram) Papineni et al.
+(2002) and ROUGE Lin (2004) for stricter quality checks.
+Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al.
+(2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020).
 A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7.
@@
 3 FOCUS
-Figure 3: Structural Causal Model (SCM) of large language models for interpreting new phrases, excluding confounders.
+This section outlines the novel approach employed in our research to enhance the adaptability of LLMs in understanding the evolving human language.
+Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
+Figure 3: Structural Causal Model (SCM) of LLMs for interpreting new phrases, excluding confounders.
 The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge.
 Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
-This section outlines the novel methodology employed in our research to enhance the adaptability of LLMs in understanding the evolving human language.
-Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
 3.1 Causal Analysis
+Figure 4: SCM analysis in FOCUS methodology.
+This figure presents the complex process of causal inference in the understanding of new phrases, highlighting how the FOCUS approach systematically analyzes and interprets intricate language patterns, emphasizing the causal links between linguistic elements and their interpretive outcomes.
 Structural Causal Models (SCMs) serve as vital tools for elucidating the relationships and influence pathways among variables.
@@
 This SCM sets the stage for our forthcoming discussion where we reintroduce and scrutinize these confounders, thus laying the groundwork for a comprehensive causal analysis.
-Figure 4: SCM analysis in FOCUS methodology.
-This figure presents the complex process of causal inference in the understanding of new phrases, highlighting how the FOCUS approach systematically analyzes and interprets intricate language patterns, emphasizing the causal links between linguistic elements and their interpretive outcomes.
-3.1.1 Causal Analysis of Entity
-Initially, we reference a typic SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context.
+Analysis of entity
+Initially, we reference a typical SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context.
 From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$.
@@
 Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
-Figure 5: The four-stage pipeline of FOCUS.
-3.1.2 Analysis with Linguistic Factors
+Analysis of linguistic factors
 In our SCM, denoted as $S=\{X,W,E,Y,R\}$, the output variable $Y$, as an endogenous variable, is influenced by the input context $X$, the input phrase $W$, other entities $E$, and finally, linguistic factors $R$, which include the linguistic structure, style, theme, and cultural background, crucially shape $X$ and constitute the exogenous variables.
@@
 This comprehensive approach allows for a nuanced understanding of how context, entities, and linguistic factors collectively influence the interpretation $Y$.
-Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
-1 2 L
-GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
-CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
-CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
-FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
-GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
-CoT Wei et al. (2022) 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
-CauView Wang et al. (2023) 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
-FOCUS (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
-Table 1: Performance results on the factual dataset.
 3.2 Context-based Causal Intervention
 Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors.
-To eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (d)).
+To concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$, (as illustrated in Figure 4 (a) and (b)).
+Further, to eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (c) and (d)).
 This alteration aids in isolating the effect of $W$ while excluding entity-specific biases.
-Further, to concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$.
-This shift accentuates the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner.
+Overall, these adjustments accentuate the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner.
 The updated causal effect formula in the SCM is thus:
@@
 The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
+Figure 5: The four-stage pipeline of FOCUS.
 The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context.
 The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language.
-To achieve this objective, we propose a four-stage method, here is a description of each stage and its principles:
+To achieve this objective, we propose a four-stage method (as shown in Figure 5), each with its principles:
 Direct Inquiry (DI)
-The DI stage is the foundational step in our FOCUS methodology.
-It primarily involves inputting the usage example (context) $X$ and the phrase $W$ directly into the language model.
-This stage is represented as $Y_{DI}=\mathcal{M}(X,W,E)$, where $\mathcal{M}$ is the function that defines the processing of the language model.
-The aim here is to assess the direct effect of the context $X$ on the inferred meaning $Y$ in a straightforward and intervention-free manner.
-This allows the model to naturally derive the meaning within the given context.
-Special emphasis is placed on maintaining the integrity of all confounders, ensuring that the phrase’s morphology, literal meaning, and context-based interpretation are coherently analyzed.
-The DI stage sets the baseline for understanding the meaning of the phrase before any further analytical layers are added in the subsequent stages.
-Masked Entity Analysis (MEA)
-The MEA stage is a crucial component of our FOCUS methodology, designed to mitigate the impact of the phrase’s literal meaning on its interpretation.
-In this stage, the phrase $W$ is masked within the context $X$, leading to an evaluation of the meaning $Y$ in the absence of direct influence from $W$.
-This process is denoted as $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, where $\mathcal{M}$ represents the language model.
-The primary objective is to understand the meaning based on context, prompting the model to propose synonyms or near-synonyms that could feasibly replace the masked phrase.
-The MEA stage allows for a nuanced analysis of how the interpretation $Y$ is influenced when $W$ is obscured, yet still impacted by the entities $E$.
-This stage is instrumental in mitigating the literal interpretation bias of $W$, focusing instead on deriving meaning from the broader linguistic factors present in $X$ and $E$.
+In the Direct Inquiry (DI) stage, we input the usage example (context) $X$ and phrase $W$ into the language model.
+Represented as $Y_{DI}=\mathcal{M}(X,W,E)$, this stage aims to evaluate the direct effect of $X$ on the inferred meaning $Y$.
+It allows the model to derive meaning naturally within the given context, focusing on the coherence of the phrase’s morphology, literal meaning, and context-based interpretation.
+DI sets a baseline for understanding the phrase’s meaning before adding more analytical layers.
+Masked Entity Inquiry (MEI)
+In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence.
+This process, $Y_{MEA},\hat{W}{MEA}=\mathcal{M}(X{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase.
+MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
+Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
+1 2 L
+GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
+CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
+CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
+FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
+GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
+CoT Wei et al. (2022) 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
+CauView Wang et al. (2023) 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
+FOCUS (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
+Table 1: Performance results on the factual dataset.
 Entity Replacement Inquiry (ERI)
-The ERI stage plays a pivotal role in our FOCUS methodology, where we explore the impact of varying entity relationships on the interpretation of phrases to delve deeper into context-based interpretation.
-In this stage, we implement an entity replacement strategy combined with a dropout rate, selectively altering entities within the context $X$.
-This process, denoted as $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$, involves generating multiple context sentences with replaced entities, while a dropout rate determines the extent of entity replacement.
-The addition of this dropout strategy is not only to introduce randomness but also to prevent the emergence of new entity biases that could arise from completely replacing all entities.
-By maintaining a mix of original and altered entities, we ensure the robustness of understanding while mitigating the risk of introducing new biases.
-The language model $\mathcal{M}$ then conducts multiple inferences on these varied contexts to synthesize a cohesive interpretation.
-This approach enhances the robustness of the model by considering a diverse range of entity interactions.
-The ERI stage is fundamental in understanding the nuanced effects of entity dynamics within different contexts, providing a more comprehensive causal analysis of phrase interpretation.
-Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
-1 2 L
-GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
-CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
-CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
-FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
-GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
-CoT Wei et al. (2022) 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
-CauView Wang et al. (2023) 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
-FOCUS (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
-Table 2: Performance results on the counterfactual dataset.
-Final Synthesis (FS)
-The FS stage is the integrative phase of the FOCUS methodology, symbolized as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$.
-This stage synthesizes insights from Direct Inquiry (DI), Masked Entity Analysis (MEA), and Entity Replacement Inquiry (ERI).
-It evaluates the interplay between direct, contextually-driven, and variable entity interpretations to forge a comprehensive understanding of the phrase $W$ within context $X$ and its relation to entities $E$.
-In FS, we assess causal relationships, reconciling the varied interpretations and confounding factors identified in previous stages.
-This synthesis refines our understanding of language model interpretations, emphasizing contextual richness and linguistic nuance.
-The FS stage thus concludes the FOCUS methodology, offering a multi-dimensional perspective on language model analytics.
+In the Entity Replacement Inquiry (ERI) stage, we alter entities in context $X$ to assess phrase interpretation variability.
+We use a dropout rate for entity alteration in $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$.
+This introduces a balance of original and new entities, enhancing model robustness without bias.
+ERI helps understand entity dynamics’ effects on interpretation, providing deeper causal analysis.
+Synthesis (SY)
+In the Synthesis (SY) stage, we integrate insights from Direct Inquiry, Masked Entity Inquiry, and Entity Replacement Inquiry.
+Represented as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$, this phase evaluates the interplay between direct, contextual, and entity-variable interpretations.
+SY reconciles varied interpretations and confounders, refining the model’s understanding of language nuances.
+This final stage offers a multi-dimensional perspective on language model analytics, emphasizing contextual richness.
 4 Experiments
@@
 We implement the CauView method in our experiment design by using a two-stage prompt inquiry, due to its lack of a direct inquiry step.
-FOCUS
-The outputs from the first three stages were processed by a parser that retained only the final conclusions, filtering out the model’s “reasoning" process.
-In the fourth stage, the language model was briefed about the concise procedures and objectives of the previous three methods.
-This briefing included potential key information each method might yield and the risks of information omission.
-Additionally, to control for syntactic variability and minimize error, the prompts for all baselines, including FOCUS, incorporated the same three sample entries for few-shot in-context learning.
-This ensured that the language model’s output syntax closely resembled that of the dataset.
-Experiment F1 Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
+4.2 Results
+Results on factual dataset
+Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases.
+Table 1 summarizes these findings.
+We can see that the FOCUS consistently outperforms other approaches across all metrics.
+Specifically, for the GPT-3.5 model, FOCUS achieves the highest $\text{F}_{1}$ score of 0.4292, precision of 0.4153, and recall of 0.4524, along with the best BLUE, ROUGE, similarity, SimCSE scores, and accuracy at 84.5%.
+A similar trend is noted with the GPT-4 model.
+Here, FOCUS again lead the metrics, notably achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an impressive accuracy of 88.2%.
+These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
+Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
+1 2 L
+GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
+CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
+CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
+FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
+GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
+CoT Wei et al. (2022) 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
+CauView Wang et al. (2023) 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
+FOCUS (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
+Table 2: Performance results on the counterfactual dataset.
+Experiment $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
 w/o MEA 0.4366 0.4380 0.4484 0.2766 0.4586 0.2059 0.3556 0.8821 0.8014 82.0
@@
 Table 3: Results of ablation experiments on GPT-4.
-4.2 Results
-Factual dataset
-Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases.
-Table 1 summarizes these findings.
-We observed that the FOCUS consistently outperformed other approaches across all metrics.
-Specifically, for the GPT-3.5 model, FOCUS achieved the highest F1 score of 0.4292, Precision of 0.4153, and Recall of 0.4524, along with the best BLUE, ROUGE, Similarity, SimCSE scores, and Accuracy at 84.5%.
-A similar trend was noted with the GPT-4 model.
-Here, FOCUS again led the metrics, notably achieving an F1 score of 0.4446, Precision of 0.4280, Recall of 0.4714, and an impressive Accuracy of 88.2%.
-These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
-The comparative analysis clearly illustrates that our FOCUS approach excels in understanding the nuanced and evolving nature of internet language, thereby marking a significant advancement in the field of natural language processing.
-Counterfactual dataset
+Results on counterfactual dataset
 Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage.
 The performance results, as detailed in Table 2, reveal insightful trends.
-Notably, our FOCUS method again excelled, outperforming others in metrics with GPT-4, achieving an F1 score of 0.4532, Precision of 0.4598, Recall of 0.4551, and an Accuracy of 84.9%.
+Notably, our FOCUS method again outperforms others in metrics with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%.
 This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios.
-Comparatively, on the factual dataset, while FOCUS led with substantial margins, the gap in performance was more pronounced in the counterfactual dataset.
-This discrepancy underscores the enhanced complexity in interpreting altered linguistic contexts, a challenge adeptly met by FOCUS.
 In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
 4.3 Ablation Study
-Our ablation study, focusing on the FOCUS methodology’s components MEA and ERI, utilized GPT-4 on the factual dataset.
+Our ablation study, focusing on the FOCUS methodology’s components MEI and ERI, utilized GPT-4 on the factual dataset.
 As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
-MEA (Masked Entity Analysis)
-MEA, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance.
-Its exclusion (w/o MEA) reduced the F1 score to 0.4366 from 0.4446.
+MEI
+The MEI stage, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance.
+Its exclusion (w/o MEI) reduced the $\text{F}_{1}$ score to 0.4366 from 0.4446.
 This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
-ERI (Entity Replacement Inquiry)
-The ERI process, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact.
-Removing ERI (w/o ERI) decreased the F1 score to 0.4283.
+ERI
+The ERI stage, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact.
+Removing ERI (w/o ERI) decreased the $\text{F}_{1}$ score to 0.4283.
 ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
-While incorporating either MEA or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results.
+While incorporating either MEI or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results.
 This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
@@
 5.1 Knowledge Update Methods
-LLMs enhance knowledge through Parameter-Efficient Fine-Tuning (PEFT) methods like task-specific parameter addition Houlsby et al.
+LLMs enhance knowledge through parameter-efficient fine-tuning methods like task-specific parameter addition Houlsby et al.
 (2019) and low-rank adaptation (LoRA) Hu et al.
@@
 Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends.
-Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new phrases and usage patterns, going beyond traditional models in precision and relevance.
-Our analysis revealed that the causal inference-based approach of FOCUS significantly outperforms traditional models in interpreting internet slang and memes.
+Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new concepts, going beyond other methods in precision and relevance.
 This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts.
@@
 This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
-7 Acknowledge
+Limitations
+The main limitation of this work lies in its reliance on Internet-specific data, primarily sourced from UrbanDictionary.
+While this provides a rich and current linguistic landscape, it may not comprehensively represent the vast diversity of global Internet slang and memes, potentially limiting the generalizability of our findings.
+Additionally, the performance of FOCUS heavily depends on the quality and variety of the dataset used.
+Our approach, although innovative, does not fully address the complexities of linguistic evolution in non-English or morphologically rich languages.
+Future work should explore more diverse datasets and extend our methodology to other languages and linguistic contexts.
+Ethics Statement
+Our research acknowledges that while methods like the SLANG benchmark and FOCUS approach enhance LLM’s understanding of Internet language, they cannot entirely eliminate the propagation of harmful content.
+Users must exercise caution and cultural sensitivity, especially when interpreting slang and memes, to avoid reinforcing stereotypes or biases.
+Our work encourages responsible use, emphasizing the importance of respecting diverse linguistic origins and the natural evolution of language.
+Acknowledge
 This paper is partially supported by the National Science Foundation of China under Grant No.U1911401, U21B2046 6237075198.

### v6

@@
 Slang: New Concept Comprehension of Large Language Models
-[…]
+Lingrui Mei1,4  Shenghua Liu1,4thanks: Corresponding author.
+Yiwei Wang2,3  Baolong Bi1,4  Xueqi Cheng1,4 1CAS Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2University of California, Los Angeles  3University of California, Merced 4University of Chinese Academy of Sciences meilingrui22@mails.ucas.ac.cn wangyw.evan@gmail.com {liushenghua,bibaolong23z,cxq}@ict.ac.cn
 1 Introduction
-Recently, the language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016).
-These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
+Recently, language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016).
+These rapid changes in language pose serious challenges to the Large Language Models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
 Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language.
 This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021).
-Hence, it is essential for LLMs to to understand linguistic shifts and new concepts without constant updates or external data.
+Hence, it is essential for LLMs to understand linguistic shifts and new concepts without constant updates or external data.
 Figure 1: Comparative analysis of LLMs’ understanding of new phrases using CoT Wei et al.
@@
 Focus significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability.
-With GPT-4, FOCUS achieved an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, and recall of 0.4551, alongside an accuracy of 84.9%, outperforming previous methods in comprehension and adaptability.
+With Claude 3, Focus achieved an $\text{F}_{1}$ score of 0.4596, precision of 0.4452, and recall of 0.4827, alongside an accuracy of 89.7%, outperforming previous methods in comprehension and adaptability.
 The codes for the Slang and Focus toolboxes are open-sourced, contributing to the community’s resources for advancing language model development.
@@
 It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability.
-We detail SLANG’s problem definition, dataset construction, and evaluation metrics in the following subsections.
-2.1 Problem
-Diving deep into the digital deluge, we encounter the fluid and fleeting world of internet parlance, where slang and idiomatic expressions defy static definitions and demand dynamic interpretation by LLMs.
-Our journey unfolds with the task of capturing the essence of this linguistic fluidity, mathematically expressed as the challenge of estimating $P(Y|W,X;\mathcal{M})$.
-This formulation seeks to unravel the likelihood of an explanation $Y$ emerging from a phrase $W$ within its context $X$, as interpreted by the model $\mathcal{M}$.
+We detail Slang’s dataset construction, and evaluation metrics in the following subsections.
 2.1 Preprocessing
@@
 Filtering
-The content filtering of our dataset consists of several steps, with the foremost being the assurance of content novelty.
-Many UrbanDictionary phrases are now common, potentially pre-included in GPT’s training data.
-Terms like ghosting or flex illustrate this shift to general usage.
-To avoid redundancy in LLM predictions, we focused on ensuring that the phrases and definitions in our dataset were new to the LLMs.
-To solve this, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models.
-To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022.
+The content filtering of our dataset consisted of several steps to ensure the quality and novelty of the concepts:
+  • Temporal Filtering: Many UrbanDictionary phrases are now common and potentially included in LLMs’ training data.
+To ensure the novelty of our dataset, we leveraged the knowledge cut-off dates of LLMs, strategically selecting phrases that emerged after these dates.
+For instance, gpt-4-0613 has a knowledge cut-off date of April 2021, as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models, and we selected concepts that were added to UrbanDictionary after January 2022.
 This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
-Another essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2.
-Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process.
-This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
-Additionally, we employed gpt-4-0613 for thorough filtration, ensuring terms were not already memorized by GPT, thus maintaining our dataset’s novelty.
-We also removed inappropriate content, preserving academic integrity and ensuring quality.
+  • User Rate Filtering: We analyzed user-generated ratings to refine the dataset.
+Entries with overwhelmingly negative receptions (more than 80% downvotes) were excluded.
+Comparative histograms in Figure 2 illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process.
+This stringent data cleaning was crucial in ensuring the quality and reliability of the entries selected for our research.
+For details on the validation of user-generated votes, see Appendix E.
+  • Removal of Inappropriate Content: Inappropriate content, including NSFW material and hate speech, was removed to preserve academic integrity and ensure quality.
+  • Novelty Check: To ensure that the concepts in our dataset were unknown to the LLMs, we employed gpt-4-0613 for thorough filtration based on the method described by Yin et al., 2023.
+Additionally, due to the existence of models with different cut-off dates, we applied the "needle in a haystack" test (Kamradt, 2023, see Appendix D) to confirm the novelty of the knowledge.
+This meticulous validation process involved strategically embedding the selected phrases into extensive corpora and evaluating the models’ capability to extract them.
+After the temporal filtering step, we started with 7220 concepts. The subsequent steps filtered out 5463, 1328, and 21 samples respectively, resulting in a final dataset of 408 usable new concepts.
 Factual dataset
-Acknowledging the informal nature of the original dataset, we transformed the data into a uniform, formal dictionary format.
-This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning.
-The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts.
-We also enriched each explanation with four synonym-based variants, ensuring the capture of diverse potential responses.
-When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches.
-This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
+Following the above filtering steps, we obtained our factual dataset.
+Acknowledging the informal nature of the original data, we transformed it into a uniform, formal dictionary format (see Appendix B).
+This process involved simplifying explanations and examples for universal understanding while preserving their original meanings.
+We adhered to a custom-designed template for consistency and clarity across all concepts.
+Each explanation was enriched with four synonym-based variants to capture diverse potential responses.
 Counterfactual Dataset
-We aim to assess the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding by preserving original meanings from the factual dataset while introducing new phrases.
-This approach maintains the semantic substance but shifts the linguistic form.
-Our strategy involves creating novel phrases that are linguistically distinct yet semantically congruent with their real-world counterparts.
-Crafting these phrases requires a creative and analytically rigorous approach, leveraging the rich contextual foundation of the original dataset.
-These new phrases encapsulate the same ideas as the original meanings but are expressed through a different linguistic perspective.
-Balancing inventive language use with logical linguistic norms is essential to ensure the new phrases are plausible and coherent.
-Upon generation, these new phrases are paired with their corresponding original meanings, creating a dataset that contrasts novel linguistic expressions against established semantic structures.
-This pairing allows for a multifaceted analysis of LLMs’ language processing and comprehension, particularly their ability to separate linguistic form from semantic substance.
-This methodology sheds light on the nuanced capabilities of LLMs in interpreting language beyond mere recognition, delving into semantics and contextual linguistics.
+To further evaluate the ability of LLMs in understanding new concepts, we created the counterfactual dataset.
+This dataset is derived from the factual dataset by preserving the original phrases while modifying the contexts and explanations.
+Given the factual dataset $\mathcal{D}_{\text{fact}}$, the generation process is structured as follows:
+  • Entity Extraction: We extract the entities $\mathbf{e}$ from each explanation $\mathbf{y}$ in $\mathcal{D}_{\text{fact}}$.
+  • Counterfactual Replacement: For each entity $e_{i}\in\mathbf{e}$, we generate a counterfactual entity $e_{i}^{\prime}$ that is conceptually divergent from the original entity $e_{i}$ while retaining the original phrase structure.
+  • Context Construction: Based on the counterfactual explanations $y_{i}^{\prime}$, we use GPT-4 to generate new contexts $x_{i}^{\prime}$.
+Consequently, the counterfactual dataset $\mathcal{D}_{\text{cf}}$ is composed of pairs: $\mathcal{D}_{\text{cf}}=\{(p_{1},x_{1}^{\prime},y_{1}^{\prime}),(p_{2},x_{2}^{\prime},y_{2}^{\prime}),\ldots,(p_{n},x_{n}^{\prime},y_{n}^{\prime})\}$.
+This approach ensures that while the original phrases are preserved, the contexts and explanations are transformed to convey entirely different meanings.
+Consequently, each entry in the counterfactual dataset introduces novel concepts that is distinct from the original dataset, providing unique challenges for the LLMs to interpret and understand.
 Figure 2: Comparative histograms illustrating the distribution of upvotes (left) and absolute upvotes (right) across dataset entries.
@@
 Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al.
-(2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020).
+(2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings of all-mpnet-base-v2 (Team, 2023; Song et al., 2020).
 A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7.
@@
 We propose a simplified SCM (see Figure 3) to delineate the interpretative processes of an LLM when confronted with novel phrases within their context.
-IIn this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation.
+In this model, users supply both the phrase $W$ and its context $X$, which are then inputted into the LLM, represented by the direct links $X\rightarrow\mathcal{M}\rightarrow Y$ and $W\rightarrow\mathcal{M}\rightarrow Y$, where $\mathcal{M}$ denotes the LLM and $Y$ is the output explanation.
 These causal links, free from confounders, capture the logical chain from input to interpretation.
@@
 Analysis of entity
-Initially, we reference a typical SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context.
+Initially, referencing a typical SCM framework (Wang et al., 2023; Wang et al., 2022), we assume SCM $S=\{X,E,Y\}$, where $X$ denotes context/input, and $E$ represents confounder, including phrases and other entities within the context.
 From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$.
@@
 The path $R\rightarrow X$ highlights the exogenous influence of linguistic factors on context.
-This configuration captures the causal mechanisms underlying natural language understanding, emphasizing the direct, indirect, and mediated relationships among the variables.
 The equation for the causal effect in this context is as follows:
@@
 $P(\tilde{W}=\tilde{w}|\tilde{E}=\tilde{e},R=r)P(\tilde{E}=\tilde{e})P(R=r)$
-This revised approach and formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention.
+This revised formula ensure a more comprehensive and nuanced analysis of the causal dynamics within the SCM, post-intervention.
 This approach ensures a more robust and bias-free interpretation within the SCM framework.
 This equation accounts for the altered relationships in the SCM after $do$-operation.
-The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context of $X$.
+The conditional probabilities and summations are now over the new variables $\tilde{E}$ and $\tilde{W}$, while maintaining the original structure’s intent to adjust for confounding effects and capture the influence of modified entities in the context $X$.
 Figure 5: The four-stage pipeline of Focus.
@@
 In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence.
-This process, $Y_{MEA},\hat{W}{MEA}=\mathcal{M}(X{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase.
+This process, $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase.
 MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
@@
 1 2 L
-GPT-3.5 Direct Ouyang et al. (2022) 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
-CoT Wei et al. (2022) 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
-CauView Wang et al. (2023) 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
-FOCUS (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
+Mistral-7B Direct Ouyang et al. (2022) 0.1869 0.2655 0.1684 0.0372 0.2078 0.0481 0.1452 0.1507 0.6084 43.1
+CoT Wei et al. (2022) 0.3453 0.3227 0.3706 0.1978 0.3657 0.1602 0.2744 0.4713 0.7132 68.5
+CauView Wang et al. (2023) 0.3012 0.2787 0.3384 0.1561 0.3272 0.1336 0.2449 0.4862 0.7127 64.3
+Focus (Ours) 0.3703 0.3555 0.3878 0.2493 0.3942 0.1866 0.3011 0.5121 0.7469 76.0
 GPT-4 Direct Ouyang et al. (2022) 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
@@
 Focus (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
-Table 1: Performance results on the factual dataset.
+Claude 3 Direct Ouyang et al. (2022) 0.2395 0.3538 0.2171 0.0552 0.2645 0.0673 0.1904 0.1947 0.6714 51.4
+CoT Wei et al. (2022) 0.4276 0.4082 0.4606 0.2471 0.4492 0.2014 0.3371 0.5628 0.7948 81.2
+CauView Wang et al. (2023) 0.3752 0.3550 0.4229 0.2015 0.4035 0.1707 0.3041 0.5778 0.7932 76.6
+Focus (Ours) 0.4596 0.4452 0.4827 0.3264 0.4835 0.2373 0.3729 0.6109 0.8354 89.7
+Table 1: Performance results on the factual dataset. For results on more models, see Appendix C.1.
 Entity Replacement Inquiry (ERI)
@@
 Large Language Models
-In our experimental setup, we utilized a subset of 408 samples, filtered from our initial dataset of 7220 concepts, focusing on evaluations over GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613).
-These models were selected for their advanced capabilities and the relevance of their knowledge bases, which have a cut-off date prior to April 2021.
-As we discussed in 2.2, this cut-off is crucial to ensure the novelty and relevancy of our dataset, as it aligns with the temporal frame of our selected UrbanDictionary concepts.
+Following the preprocessing method described in Section 2.1, we filtered our initial dataset of 7220 concepts, resulting in 408 new concepts.
+These evaluations were conducted using Claude 3, GPT-4, Mistral-7B, and other popular models.
+For detailed experimental setup and additional results for other models, please refer to Appendix A and Appendix C.
 Baselines
@@
 4.2 Results
-Results on factual dataset
-Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases.
-Table 1 summarizes these findings.
-We can see that the FOCUS consistently outperforms other approaches across all metrics.
-Specifically, for the GPT-3.5 model, FOCUS achieves the highest $\text{F}_{1}$ score of 0.4292, precision of 0.4153, and recall of 0.4524, along with the best BLUE, ROUGE, similarity, SimCSE scores, and accuracy at 84.5%.
-A similar trend is noted with the GPT-4 model.
-Here, FOCUS again lead the metrics, notably achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an impressive accuracy of 88.2%.
-These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
+4.3 Experimental Results
+Results on the factual dataset
+Table 1 summarizes the findings that the Focus method demonstrates exceptional performance across all models.
+For GPT-4, Focus achieves the highest $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an accuracy of 88.2%.
+Similarly, Claude 3 under Focus secures an $\text{F}_{1}$ score of 0.4596 and an accuracy of 89.7%, while Mistral 7B records an $\text{F}_{1}$ score of 0.3703 with an accuracy of 76.0%.
+These metrics highlight the effectiveness of the Focus method in enhancing the interpretative capabilities of language models.
 Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
 1 2 L
-GPT-3.5 Direct Ouyang et al. (2022) 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
-CoT Wei et al. (2022) 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
-CauView Wang et al. (2023) 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
-FOCUS (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
+Mistral-7B Direct Ouyang et al. (2022) 0.1765 0.2402 0.1461 0.0327 0.1983 0.0396 0.1378 0.1520 0.5488 20.0
+CoT Wei et al. (2022) 0.3318 0.3018 0.3741 0.1849 0.3593 0.1594 0.2734 0.3960 0.6692 56.2
+CauView Wang et al. (2023) 0.2922 0.2567 0.3416 0.1431 0.3197 0.1233 0.2365 0.3515 0.6309 46.3
+Focus (Ours) 0.3935 0.3928 0.4005 0.2498 0.4151 0.1891 0.3199 0.5079 0.7343 77.5
 GPT-4 Direct Ouyang et al. (2022) 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
@@
 Focus (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
-Table 2: Performance results on the counterfactual dataset.
+Claude 3 Direct Ouyang et al. (2022) 0.2123 0.3143 0.1779 0.0454 0.2435 0.0496 0.1769 0.1733 0.5909 23.1
+CoT Wei et al. (2022) 0.3928 0.3605 0.4405 0.2143 0.4138 0.1867 0.3184 0.4499 0.7398 64.7
+CauView Wang et al. (2023) 0.3468 0.3101 0.4062 0.1662 0.3791 0.1490 0.2879 0.4012 0.6941 52.4
+Focus (Ours) 0.4636 0.4739 0.4618 0.3132 0.4894 0.2361 0.3867 0.5783 0.8216 86.8
+Table 2: Performance results on the counterfactual dataset. For results on more models, see Appendix C.2.
 Experiment $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
@@
 w/o ERI 0.4283 0.4300 0.4371 0.2814 0.4593 0.2117 0.3547 0.9021 0.8092 84.0
-Table 3: Results of ablation experiments on GPT-4.
-Results on counterfactual dataset
+Table 3: Results of ablation experiments on GPT-4. For results on more models, see Appendix C.3.
+Results on the counterfactual dataset
 Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage.
-The performance results, as detailed in Table 2, reveal insightful trends.
-Notably, our FOCUS method again outperforms others in metrics with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%.
-This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios.
-In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
+As detailed in Table 2, the Focus method outshines other techniques, particularly with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%.
+Claude 3 also performs well under Focus, securing an $\text{F}_{1}$ score of 0.4636 and an accuracy of 86.8%.
+Mistral 7B shows a solid performance, with an $\text{F}_{1}$ score of 0.3935 and an accuracy of 77.5%.
+These results underscore Focus’s robust ability to navigate the challenges posed by modified linguistic contexts.
 4.4 Ablation Study
-Our ablation study, focusing on the FOCUS methodology’s components MEI and ERI, utilized GPT-4 on the factual dataset.
+Our ablation study, focusing on the Focus methodology’s components MEI and ERI on the factual dataset.
 As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
+Extended ablation results are provided in Appendix C.3.
 MEI
@@
 (2020).
-Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al.
+However, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al.
 (2021).
 BitFit Ben-Zaken et al.
-(2021), while simplifying fine-tuning, relies heavily on dataset quality.
-Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods.
+(2021) simplifies fine-tuning but relies heavily on dataset quality.
+Trade-offs in computational demands, flexibility, and task compatibility are essential considerations.
 Retrieval-augmented generation Lewis et al.
@@
 Counterfactual methods for debiasing Chen et al.
-(2023) and eliminating shortcuts Wen et al.
+(2023b) and eliminating shortcuts Wen et al.
 (2022) have shown promise, though limitations remain in targeting white-box models and retraining requirements.
-Our research addresses these limitations through contextual learning and causal inference for bias and shortcut elimination without retraining.
 6 Conclusion
-In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs).
+In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of LLMs.
 Our study introduced a novel benchmark, Slang, to assess LLMs’ proficiency in comprehending emerging linguistic trends.
-Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new concepts, going beyond other methods in precision and relevance.
-This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts.
+Additionally, we proposed the Focus methodology, which utilizes causal inference to enhance understanding of new concepts, going beyond other methods in terms of precision and relevance.
+Our approach involves the construction of datasets from UrbanDictionary, a platform known for user-generated content that reflects current language trends.
+We incorporated both factual and counterfactual instances to provide diverse linguistic contexts.
+Factual instances are drawn directly from the UrbanDictionary entries, while counterfactual instances are created by altering real-world examples to assess the models’ adaptability to hypothetical scenarios.
 The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our Focus method, to adapt to the rapid evolution of online language.
-This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
+This research contributes to the field of natural language processing by emphasizing the importance of contextual understanding and adaptability in LLMs.
+Our findings suggest that LLMs can effectively navigate the complexities of evolving human communication when equipped with robust methodologies like Focus and evaluated against benchmarks such as Slang.
 Limitations
-The main limitation of this work lies in its reliance on Internet-specific data, primarily sourced from UrbanDictionary.
-While this provides a rich and current linguistic landscape, it may not comprehensively represent the vast diversity of global Internet slang and memes, potentially limiting the generalizability of our findings.
-Additionally, the performance of FOCUS heavily depends on the quality and variety of the dataset used.
-Our approach, although innovative, does not fully address the complexities of linguistic evolution in non-English or morphologically rich languages.
-Future work should explore more diverse datasets and extend our methodology to other languages and linguistic contexts.
+The main limitation of this work is that it does not fully address the complexities of linguistic evolution in non-English or morphologically rich languages.
+Therefore, future work should explore a wider range of linguistic scenarios and extend our methodology to other languages and linguistic contexts.
+Additionally, the Focus methodology, despite its effectiveness in enhancing the understanding of emerging linguistic phenomena, has a higher computational complexity compared to some traditional approaches.
+This might not only increase the computational demands but also introduce delays when deployed on mobile devices, which could hinder real-time applications.
+Such issues necessitate further optimization to reduce computational load and improve efficiency for mobile and other constrained environments.
+Moreover, applying this methodology to downstream tasks might encounter challenges related to data processing, as the sources of new concepts may not be readily accessible.
+This could require the use of external tools to gather relevant data, potentially making the data collection process time-consuming and variable depending on the specific task.
 Ethics Statement
@@
 Acknowledge
-This paper is partially supported by the National Science Foundation of China under Grant No.U1911401, U21B2046 6237075198.
+This paper is partially supported by the National Science Foundation of China under Grant No.U21B2046 and 6237075198, and National Key R&D Program of China (No.2023YFC3305303).
+Appendix A Detail Experimental Setup
+In this section, we provide a comprehensive overview of our experimental setup, including the models used, API parameters, preprocessing steps, and the standardization process for definitions.
+This setup ensures that our evaluations are thorough, replicable, and provide meaningful insights into the performance of various large language models (LLMs).
+A.1 Models
+We evaluated both closed-source and open-source LLMs to ensure a broad assessment.
+Closed-source models.
+  • GPT-3.5:
+    gpt-3.5-turbo-1106
+  • GPT-4:
+    gpt-4-0613
+  • Claude 3 Opus:
+    claude-3-opus-20240229
+  • Claude 3 Sonnet:
+    claude-3-sonnet-20240229
+  • Claude 3 Haiku:
+    claude-3-haiku-20240307
+Open-source models.
+  • Mistral 7B Jiang et al. (2023) :
+    mistral-7b-instruct-v0.1
+  • LLaMA 7B:
+    llama-2-7b-chat
+A.2 API Parameters
+To ensure consistency and comparability in our evaluations, we used the following hyperparameters when calling the LLM API:
+  • Temperature: 0.7
+  • Max tokens: 512
+A.3 Preprocessing
+The selection of 408 samples from 7220 was based strictly on the data filtering strategy described in Section 2.1 of our paper.
+Although this number of test samples may seem small for an unfiltered dataset, it is substantial compared to existing LLM benchmarks:
+  • MaliciousInstruct (Huang et al., 2023):
+    100 samples
+  • HumanEval (Chen et al., 2021):
+    163 samples
+  • AdvBench (Zou et al., 2023):
+    500 samples
+  • HarmBench (Mazeika et al., 2024):
+    400 unimodal samples
+Additionally, we ensured the quality and diversity of the samples, as illustrated in Figure 6.
+Our chosen test samples were novel to the LLM, guaranteeing a comprehensive evaluation of model performance by maintaining the integrity and representativeness of the dataset.
+A.4 Costs
+The experiments conducted using closed-source models incurred a total cost of approximately \$500 (includes API discounts).
+For the open-source models, the experiments were run on a server with 4 NVIDIA Tesla A100 GPUs for a duration of 14 hours.
+This breakdown of costs highlights the computational and financial resources required to conduct comprehensive evaluations of large language models.
+The use of both closed-source and open-source models ensures a diverse and robust assessment, while the detailed cost analysis provides transparency regarding the experimental setup.
+Appendix B Explanation Standardization
+Since the definitions were user-generated, they varied widely in language style and structural format. To ensure uniformity, we standardized these explanations using a consistent template.
+We used the following template to standardize the explanations in the dataset:
+[P] refers to [B]. It is often used [C]. This expression [A].
+  • P: Phrase
+  • B: Basic description of the word
+  • C: Context or situation of usage
+  • A: Additional details like connotations, emotions, or typical reactions associated with the word
+For example, the original explanation of the phrase The Winter Arc was:
+“A time where every man must face the mental and physical challenges of winter. A time to put your head down and get things done”
+After standardization, it became:
+“The Winter Arc refers to a time when people deal with the cold and hard parts of winter.
+It is often used to talk about staying strong and getting work done even when it’s cold and challenging outside.
+This expression suggests that people are being tough and focused.”
+Appendix C Additional Experimental Results
+We provide a comprehensive summary of the extended evaluation conducted on various language models including GPT-3.5, GPT-4, several versions of Claude 3, Mistral 7B, and LLaMA 2-7B Touvron et al.
+(2023).
+Each model undergoes assessment using a range of prompting methods, such as Direct, CoT, CauView, and our Focus method.
+C.1 Factual Dataset
+In the factual dataset, the Focus prompting method propels GPT-4 and Claude 3 Opus models to the highest performance, with GPT-4 achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and accuracy of 88.2%.
+Claude 3 Opus closely follows with an $\text{F}_{1}$ score of 0.4596 and accuracy of 89.7%.
+The LLaMA 2-7B also exhibits commendable improvements, confirming the efficacy of Focus across diverse architectures.
+These results are shown in Table 4.
+Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLEU ROUGE Similarity SimCSE ACC (%)
+1 2 L
+GPT-3.5 Direct 0.2219 0.1541 0.4320 0.0477 0.2410 0.0547 0.1531 0.1943 0.6806 47.6
+CoT 0.3922 0.3585 0.4583 0.2273 0.4181 0.1822 0.3135 0.5292 0.7653 76.4
+CauView 0.3500 0.2487 0.4885 0.1676 0.3777 0.1506 0.2676 0.5612 0.7868 72.0
+Focus (Ours) 0.4292 0.4153 0.4524 0.2798 0.4541 0.2131 0.3481 0.5748 0.8017 84.5
+GPT-4 Direct 0.2308 0.3474 0.1917 0.0483 0.2597 0.0606 0.1859 0.1616 0.6476 47.2
+CoT 0.4123 0.3947 0.4244 0.2384 0.4370 0.1927 0.3299 0.5521 0.7883 79.3
+CauView 0.3602 0.3444 0.3948 0.1987 0.3917 0.1643 0.3032 0.5515 0.7636 74.7
+Focus (Ours) 0.4446 0.4280 0.4714 0.3177 0.4721 0.2332 0.3652 0.6032 0.8216 88.2
+Claude 3 Opus Direct 0.2395 0.3538 0.2171 0.0552 0.2645 0.0673 0.1904 0.1947 0.6714 51.4
+CoT 0.4276 0.4082 0.4606 0.2471 0.4492 0.2014 0.3371 0.5628 0.7948 81.2
+CauView 0.3752 0.3550 0.4229 0.2015 0.4035 0.1707 0.3041 0.5778 0.7932 76.6
+Focus (Ours) 0.4596 0.4452 0.4827 0.3264 0.4835 0.2373 0.3729 0.6109 0.8354 89.7
+Claude 3 Sonnet Direct 0.2251 0.3411 0.2015 0.0495 0.2541 0.0592 0.1813 0.1587 0.6417 46.8
+CoT 0.4075 0.3895 0.4202 0.2353 0.4312 0.1892 0.3247 0.5467 0.7834 78.7
+CauView 0.3562 0.3393 0.3897 0.1946 0.3872 0.1602 0.2951 0.5462 0.7586 73.9
+Focus (Ours) 0.4391 0.4238 0.4653 0.3128 0.4676 0.2293 0.3601 0.5982 0.8167 87.6
+Claude 3 Haiku Direct 0.2137 0.3105 0.1912 0.0463 0.2356 0.0558 0.1652 0.1849 0.6639 48.7
+CoT 0.3853 0.3530 0.4271 0.2196 0.4085 0.1776 0.3050 0.5217 0.7559 75.2
+CauView 0.3384 0.3032 0.3816 0.1782 0.3663 0.1498 0.2743 0.5394 0.7562 70.6
+Focus (Ours) 0.4086 0.3875 0.4308 0.2759 0.4326 0.2041 0.3303 0.5648 0.7896 82.8
+Mistral-7B Direct 0.1869 0.2655 0.1684 0.0372 0.2078 0.0481 0.1452 0.1507 0.6084 43.1
+CoT 0.3453 0.3227 0.3706 0.1978 0.3657 0.1602 0.2744 0.4713 0.7132 68.5
+CauView 0.3012 0.2787 0.3384 0.1561 0.3272 0.1336 0.2449 0.4862 0.7127 64.3
+Focus (Ours) 0.3703 0.3555 0.3878 0.2493 0.3942 0.1866 0.3011 0.5121 0.7469 76.0
+LLaMA 2-7B Direct 0.1581 0.2115 0.1361 0.0282 0.1795 0.0378 0.1254 0.1134 0.5792 40.5
+CoT 0.3174 0.2884 0.3491 0.1746 0.3344 0.1407 0.2515 0.4628 0.6844 64.4
+CauView 0.2717 0.2441 0.3109 0.1416 0.2951 0.1152 0.2242 0.4669 0.6718 59.5
+Focus (Ours) 0.3391 0.3282 0.3549 0.2043 0.3619 0.1579 0.2771 0.4873 0.7117 70.8
+Table 4: Performance results on the factual dataset.
+C.2 Counterfactual Dataset
+When analyzing the counterfactual dataset, which involves challenges from hypothetical language alterations, Focus maintains the lead under GPT-4, achieving an $\text{F}_{1}$ score of 0.4532 and accuracy of 84.9%.
+Claude 3 Opus exhibits robustness in this modified context with an $\text{F}_{1}$ score of 0.4636 and accuracy of 86.8%.
+The lower-resource models such as LLaMA 2-7B also displays significant gains, demonstrating the adaptability of Focus to a wide range of models and scenarios.
+These results are shown in Table 5.
+Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLEU ROUGE Similarity SimCSE ACC (%)
+1 2 L
+GPT-3.5 Direct 0.1922 0.1252 0.4497 0.0371 0.2120 0.0477 0.1328 0.1936 0.6051 24.3
+CoT 0.3576 0.3138 0.4407 0.2098 0.3865 0.1713 0.2932 0.4390 0.7232 62.3
+CauView 0.3161 0.2489 0.4471 0.1464 0.3439 0.1297 0.2491 0.4031 0.6954 54.6
+Focus (Ours) 0.4078 0.4018 0.4294 0.2519 0.4344 0.1955 0.3339 0.5594 0.7836 83.4
+GPT-4 Direct 0.2050 0.3018 0.1705 0.0426 0.2341 0.0503 0.1693 0.1645 0.5735 21.0
+CoT 0.3821 0.3573 0.4247 0.2189 0.4091 0.1841 0.3147 0.4383 0.7241 62.6
+CauView 0.3357 0.3024 0.3919 0.1744 0.3701 0.1467 0.2812 0.3815 0.6715 47.2
+Focus (Ours) 0.4532 0.4598 0.4551 0.3017 0.4763 0.2273 0.3722 0.5645 0.8065 84.9
+Claude 3 Opus Direct 0.2123 0.3143 0.1779 0.0454 0.2435 0.0496 0.1769 0.1733 0.5909 23.1
+CoT 0.3928 0.3605 0.4405 0.2143 0.4138 0.1867 0.3184 0.4499 0.7398 64.7
+CauView 0.3468 0.3101 0.4062 0.1662 0.3791 0.1490 0.2879 0.4012 0.6941 52.4
+Focus (Ours) 0.4636 0.4739 0.4618 0.3132 0.4894 0.2361 0.3867 0.5783 0.8216 86.8
+Claude 3 Sonnet Direct 0.1984 0.2886 0.1637 0.0398 0.2261 0.0445 0.1628 0.1539 0.5540 20.5
+CoT 0.3752 0.3439 0.4211 0.2038 0.3972 0.1790 0.3053 0.4249 0.7080 61.1
+CauView 0.3302 0.2952 0.3873 0.1578 0.3617 0.1421 0.2745 0.3776 0.6596 49.6
+Focus (Ours) 0.4416 0.4521 0.4398 0.2956 0.4666 0.2249 0.3682 0.5502 0.7874 83.2
+Claude 3 Haiku Direct 0.1853 0.2392 0.1482 0.0334 0.2018 0.0384 0.1405 0.1708 0.6026 22.7
+CoT 0.3458 0.3044 0.3898 0.1923 0.3649 0.1589 0.2763 0.4297 0.7196 60.5
+CauView 0.3038 0.2583 0.3540 0.1490 0.3280 0.1234 0.2431 0.3860 0.6786 51.7
+Focus (Ours) 0.3959 0.3885 0.4086 0.2462 0.4164 0.1868 0.3180 0.5428 0.7716 81.3
+Mistral 7B Direct 0.1765 0.2402 0.1461 0.0327 0.1983 0.0396 0.1378 0.1520 0.5488 20.0
+CoT 0.3318 0.3018 0.3741 0.1849 0.3593 0.1594 0.2734 0.3960 0.6692 56.2
+CauView 0.2922 0.2567 0.3416 0.1431 0.3197 0.1233 0.2365 0.3515 0.6309 46.3
+Focus (Ours) 0.3935 0.3928 0.4005 0.2498 0.4151 0.1891 0.3199 0.5079 0.7343 77.5
+LLaMA 2-7B Direct 0.1421 0.1680 0.1107 0.0207 0.1495 0.0281 0.1021 0.1033 0.4606 13.7
+CoT 0.2619 0.2381 0.2950 0.1472 0.2837 0.1255 0.2159 0.3151 0.5475 44.5
+CauView 0.2309 0.2026 0.2698 0.1122 0.2520 0.0971 0.1864 0.2762 0.5177 36.1
+Focus (Ours) 0.3087 0.3081 0.3142 0.1954 0.3255 0.1479 0.2510 0.3985 0.6073 60.6
+Table 5: Performance results on the counterfactual dataset.
+C.3 Ablation Study
+We extend the ablation study to more models, focusing on the effect of the MEI and ERI components of Focus on the factual dataset.
+As shown in Table 6, removing these components leads to a significant performance drop across all models tested.
+These results further confirm the findings in Section 4.4, demonstrating that MEI and ERI are crucial for achieving optimal performance within Focus.
+C.4 Discussion
+Across both datasets, Focus consistently excels, enabling models to understand and interpret language effectively.
+This effectiveness is particularly evident in models with diverse capacities and structures, signifying the versatility of Focus in its application to natural language processing tasks.
+Furthermore, the lower recall score observed in GPT-3.5 is attributed to the model’s tendency to provide shorter answers to questions, a characteristic preference of GPT-3.5 itself rather than a limitation of Focus.
+Although this preference for brevity contributes to a lower recall score, it does not diminish the overall effectiveness of Focus.
+Model Experiment $\mathbf{F}_{1}$ Precision Recall BLEU ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC(%)
+Claude 3 Opus w/o MEA 0.4325 0.4160 0.4561 0.2962 0.4565 0.2153 0.3480 0.5806 0.8018 85.1
+w/o ERI 0.4369 0.4202 0.4608 0.3003 0.4606 0.2184 0.3519 0.5857 0.8062 85.8
+GPT-4 MEA 0.4366 0.4380 0.4484 0.2766 0.4586 0.2059 0.3556 0.8821 0.8014 82.0
+w/o ERI 0.4283 0.4300 0.4371 0.2814 0.4593 0.2117 0.3547 0.9021 0.8092 84.0
+Mistral 7B w/o MEA 0.3624 0.3477 0.3792 0.2414 0.3864 0.1787 0.2933 0.5032 0.7379 74.7
+w/o ERI 0.3663 0.3515 0.3833 0.2453 0.3903 0.1816 0.2972 0.5081 0.7421 75.4
+Table 6: Results of ablation experiments on more models.
+Appendix D Needle In A Haystack Test
+To ensure that the concepts in our dataset are novel to LLMs, we designed and implemented the "needle in a haystack" test.
+The purpose of this test is to evaluate the ability of LLMs to retrieve specific information from large corpora, thereby verifying whether the concepts we selected are indeed new to the LLMs.
+D.1 Experimental Setup
+A large corpus was prepared, incorporating text data from diverse online sources such as news articles, blog posts, and social media content.
+This corpus is intended to simulate the training data typically encountered by LLMs.
+We selected a set of concepts that were added to our Urban Dictionary dataset after the LLMs’ knowledge cutoff date.
+For each concept, we generated a unique phrase embedding the concept and randomly inserted this phrase into the corpus.
+This setup was designed to mimic the challenge of locating specific information in a vast dataset.
+We tested the LLMs’ ability to identify and extract each phrase from the corpus using specific prompts.
+If an LLM successfully retrieved the inserted phrase, the corresponding concept was considered known to the model.
+If the model failed to find the phrase, the concept was deemed novel.
+Figure 6: Distribution of internet slang categories across our dataset.
+Volunteers 1 2 3 4 5 Average
+Consistency (%) 96.50 97.00 96.00 97.50 98.00 97.00
+Table 7: Consistency of the user votes-based filtering strategy across different volunteers.
+Volunteers Recall (%) Precision (%)
+Volunteer 1 95.00 94.12
+Volunteer 2 97.00 93.27
+Volunteer 3 95.00 95.00
+Volunteer 4 98.00 92.45
+Volunteer 5 99.00 91.67
+Average 96.80 93.30
+Table 8: Recall and precision of the user votes-based filtering strategy as validated by human reviewers.
+D.2 Data Format Example
+Here is an illustrative example of the data format used in the "needle in a haystack" test:
+  • Concept: "Tamagotchi effect"
+  • Inserted Phrase: "Jimmy was so upset when his furby died, he obviously was suffering from the tamagotchi effect."
+  • Corpus Sample: "…The stock market showed surprising resilience today.
+In other news, Jimmy was so upset when his furby died, he obviously was suffering from the tamagotchi effect.
+Meanwhile, local sports teams are gearing up for the upcoming championships…"
+  • LLM Prompt (w/o few-shot context): "…Identify the phrase from the text that describes a scenario where a person shows emotional distress due to the cessation of function in an electronic device or machine…"
+  • LLM Output: "Jimmy was so upset when his furby died, he obviously was suffering from the tamagotchi effect."
+If the LLM accurately extracts the inserted phrase, "tamagotchi effect" would be considered known by the LLM.
+If not, it is marked as a novel concept.
+This process was repeated for all selected concepts to determine their novelty to the LLMs.
+D.3 Discussion
+It is also important to note that in the context of this experiment, the use of existing filtering strategies based on the cut-off date for GPT models proved to be more stringent than the "needle in a haystack" test.
+This resulted in no additional data being filtered out by the test since the dataset had already been screened through more conservative criteria Yin et al.
+(2023) and aligned with an earlier knowledge cut-off date.
+However, this situation is specific to the dataset used in this study, which is a subset tailored to demonstrate the methodology.
+For broader applications, especially when utilizing our complete open-source dataset which contains over 180,000 entries, the "needle in a haystack" test becomes essential.
+This is crucial for effectively assessing the novelty of concepts across a more extensive and diverse corpus.
+Appendix E Validation of User-generated Votes
+To validate the effectiveness of our user votes-based filtering strategy, we conducted an experiment with the assistance of five English-speaking volunteers from English-speaking countries.
+This experiment was designed to compare the human judgment against the automated user votes-based method, affirming the reliability of user votes as a metric for assessing data quality.
+E.1 Experiment Design
+The experiment engaged five volunteers to review 200 entries from our dataset.
+These entries included an equal split of 100 entries that had been filtered out and 100 that had been retained by our pre-existing user votes-based filtering method.
+Volunteers were instructed to make independent filtering decisions for each entry based on our study’s quality criteria.
+E.2 Data Collection and Analysis
+We analyzed the decisions of the volunteers to determine the recall and precision of the user votes method against human judgments.
+Recall measures the proportion of entries that both humans and the automated method agreed should be filtered, while precision assesses the accuracy of the automated method in filtering entries deemed necessary by human reviewers.
+Additionally, we calculated consistency rates to quantify the agreement between each volunteer’s decisions and the automated method.
+E.3 Results
+The data, as shown in Table 8 and 7, indicate high recall rates (95% to 99%) and precision rates (91.67% to 94.12%), along with very high consistency rates (96.50% to 98.00%).
+These metrics collectively demonstrate that the user votes-based method is highly effective at mirroring human judgment in filtering decisions.
+The results underscore the potential of user votes as a reliable indicator of content quality, validating its use as a principal method for data filtering in our study.
+Appendix F Dataset Categorization
+To address potential biases in our dataset that could arise from overrepresentation of certain internet slang categories, such as metaphors, we classified the slang phrases into eight broad categories.
+We recruited five English-speaking volunteers from English-speaking countries to assist in the categorization of 408 new concepts used in our experiments.
+Each volunteer, drawing on their personal experience and familiarity with internet culture, independently categorized each phrase. The categories were as follows:
+  • Abbreviations and acronyms
+  • Pop culture references
+  • Technical and internet terms
+  • Metaphors and similes
+  • Gaming and subculture jargon
+  • Euphemisms and slang for sensitive topics
+  • Social media and communication shortcuts
+  • Generational and temporal slang
+This collaborative approach was crucial to ensuring that our dataset was not skewed toward any single category of slang, providing a more balanced foundation for analysis.
+As shown in Figure 6, our dataset features a diverse array of slang expressions.
+Concept Editing and Tuning
+Recent advancements in concept editing and tuning have significantly improved the problem-solving capabilities of models across a wide range of fields (Zhang et al., 2023a, 2024b, 2024c; Li et al., 2024; Chen et al., 2023a; Zhang et al., 2024a).
+These methods, which modify the internal structure of large language models (LLMs), are designed to adjust the output based on newly edited knowledge.
+In particular, many techniques focus on integrating auxiliary networks or tweaking model parameters to guide responses (Meng et al., 2022a, b; Mitchell et al., 2022; Yao et al., 2023; Bi et al., 2024c).
+A promising approach in this area is In-Context Editing (ICE)(Bi et al., 2024e, a, b, d), which enables models to adapt by utilizing prompts with modified facts and retrieving relevant editing demonstrations from memory.
+However, hallucinations and safety issues remain significant challenges in tasks related to LLMs(Zhang et al., 2023b; Mei et al., 2024a, b).
+Appendix G Summary of Contributions
+While our experiments leverage the UrbanDictionary dataset and focus on a subset of popular LLMs, it is crucial to emphasize that the core contributions of this work extend far beyond these specifics.
+We would like to highlight three key aspects that underscore the broader impact and applicability of our research:
+Dataset construction pipeline
+This work goes beyond simply providing a static dataset from a single source.
+Instead, we have developed a comprehensive, open-source toolbox that empowers researchers and practitioners to continuously collect, clean, and process data from a wide range of sources.
+This toolbox is designed to be highly adaptable, allowing users to easily integrate and analyze data from various platforms and domains, such as social media, online forums, and digital publications.
+By offering a flexible and extensible framework, our approach ensures the long-term relevance and applicability of the methodology, enabling researchers to keep pace with the ever-evolving landscape of online language.
+The open-source nature of the toolbox further encourages collaboration and innovation within the research community, as it allows anyone to leverage and build upon our work to process and analyze data from diverse sources, tailoring it to their specific research questions and requirements.
+Benchmarking framework
+The Slang benchmark is not merely a one-off evaluation limited to the specific datasets and models used in our experiments.
+Rather, it presents a comprehensive and generalized framework for assessing the adaptability and comprehension capabilities of a wide range of large language models when faced with the challenges of evolving linguistic phenomena.
+The benchmark is designed to be model-agnostic and can be seamlessly applied to various datasets and architectures, irrespective of their size, domain, or underlying structure, allowing for standardized evaluations and fair comparisons across different settings.
+By providing a robust and flexible evaluation framework, Slang sets a new standard for assessing the performance of large language models in the face of linguistic change, facilitating the development of language models that can handle the dynamic nature of human language and paving the way for more adaptable and resilient natural language processing systems.
+Enhancing LLMs on the fly
+The Focus methodology proposed in this work offers a principled approach for improving the ability of LLMs to grasp, interpret, and adapt to emerging linguistic phenomena on the fly, without the need for retraining or relying on Retrieval-augmented generation (RAG) techniques.
+Our method is model-agnostic, providing a strong basis for anticipating its applicability and benefits across a broader spectrum of LLMs and architectures, without the computational overhead and data requirements associated with retraining or RAG-based approaches.
+This positions Focus as an efficient and scalable solution for enhancing LLMs’ understanding of new concepts.
+In industrial applications, Focus has the potential to empower businesses across various domains, enabling real-time product recommendation systems, content moderation, sentiment analysis tools, and customer service chatbots that can swiftly adapt to emerging new concepts and terminology.
+By enhancing enterprises’ responsiveness and adaptability, Focus positions itself as a cost-effective solution that can provide a significant competitive advantage in the fast-paced digital market.
+Appendix H Definitions
+Concept
+refers to new ideas or phenomena emerging in language due to human activities, particularly on the internet.
+Expression
+is a specific phrase or term used to convey these concepts.
+Deeper meaning
+refers to the underlying significance or implications of a concept beyond its literal expression.
+Linguistic shift
+denotes the gradual incorporation of new concepts into the language, leading to changes over time.